OpenStudy (anonymous):

What is the sum or difference? 3x^8-7x^8

OpenStudy (raden):

3x^8-7x^8 = (3-7)x^8 = -4x^8

OpenStudy (anonymous):

6y^5 – 9y^5

OpenStudy (raden):

like the first, just subtract the coefficients each term y^5, 6y^5 – 9y^5 = (6-9)y^5 = -3y^5